# thanks: These authors contributed equallythanks: These authors contributed equallythanks: These authors contributed equallythanks: Present address: Nord Quantique, Sherbrooke, QC J1J 2E2, Canada

### v1

@@ IV Relevance for chemistry
Figure 6: Lindbladian simulation of an ordinary double-well system. A) Dynamics simulated as a function of well-asymmetry $k_{1}/k_{4}$ and well-depth $k_{2}/k_{4}$. The color code marks the activation time (compare to Fig. 4 B). The green-black line shows the maximum activation time. B) Comparison of the activation time for the ordinary symmetric ($k_{1}=0$) double-well and the (optimal) activation time along the green-black line in A as a function of well-depth $k_{2}/k_{4}$. C) Activation time as a function of asymmetry for $k_{2}/k_{4}=12.6$ (black dashed line in A), showing resonant tunneling (as in Fig. 3, see also [5]). Note that, as in the parametric oscillator Fig. 3, the linewidths alternate from broad to narrow.
The experimental observations presented here raise the question of whether they are present in other double-well systems, like the type of double-well used to describe chemical reactions. This may prove useful for investigations of problems relevant to chemistry (see, for example, [5] and references therein). To investigate this, we run a Lindbladian simulation for the chemical system with Hamiltonian $\hat{H}/\hbar=\hat{p}^{2}/2+k_{4}\hat{x}^{4}-k_{2}\hat{x}^{2}+k_{1}\hat{x}$ and including single photon loss and gain with rate $\kappa/k_{4}=0.025$ and at temperature $n_{\mathrm{th}}=0.05$ (parameters that correspond to the ranges accessible by our system). Note that the dissipation model is still that of a high-quality factor linear oscillator, and we work under the hypothesis that this dissipation model captures the relevant phenomenology for applications in chemistry. With this, we answer our question affirmatively. On the one hand, the green-black line in Fig. 6 A depicts the maximum lifetime as a function of well-asymmetry and well-depth, showing that - just like in our experiment (Fig. 4 A, B) - a small asymmetry increases the lifetime significantly (see also Fig. 5 A, B and Fig. 6 B). On the other hand, Fig. 6 C depicts a horizontal linecut of Fig. 6 A and clearly reveals the same width alternation (broad-narrow-broad) as discovered and explained in our experiment (see Fig. 3). The finding of these two unexpected effects shows that our asymmetric Kerr parametric oscillator setup is already able to produce meaningful predictions for chemical quantum rate theory.
Double-well systems have been extensively studied in various contexts [32, 33], including in superconducting circuits [34, 35, 36, 37]. However, we speculate that the unique combination of precise real-time microwave control, complete tunability over Hamiltonian parameters, experimental stability, fast repetition rates, and high-fidelity readout in our setup is the reason why we observed effects not previously reported.
@@ V Conclusion
We also note the importance of this system for quantum computation since qubits can be encoded in the well state manifold [22, 18]. In this regard, two contributions of the present work deserve to be highlighted. The first one is the increase of the activation timescale $T$ by a fine control of the asymmetry. This directly leads to a reduction of bit-flip errors [42] with no extra hardware requirements. The second one is that we demonstrated the operation of a highly asymmetric parametric oscillator in the quantum regime. We provide direct evidence that the static effective description is not compromised under strong linear drives, which are required for fast gates [22] and new implementations of hardware efficient readout schemes [43].
Our control over the well-asymmetry opens the path for quantum simulation of chemical reaction dynamics with a parametric oscillator [38]. The herein shown controllable asymmetry and well depth, in conjunction with adjustable detunings [24], allows for the emulation of chemical double well potential energy surfaces [38]. This in turn allows for the analog simulation of proton tunneling, and e.g. the study of transfer reactions between the Guanine-Cytosine DNA base pairs appears within reach of current Kerr parametric oscillator technology [38].
@@
ACCdA, RGC, and MS analyzed the data, modeled the experiment, and wrote the manuscript. PEV supported the data analysis. All authors contributed to the manuscript. RGC and MS conceptualized the experiment and collected data. NEF fabricated the device. BA, DGAC, PK, EG, and VSB validated the relevance and utility to chemistry and provided theory support. MHD and RGC supervised the research.
@@ Acknowledgements
We acknowledge discussions with Q. Su, A. Ding, B. Brock and M. Dykman. This research was sponsored by the Army Research Office (ARO) under award number W911NF-23-1-0051. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office (ARO), or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. VSB, EG and MHD acknowledge partial support from the National Science Foundation Center for Quantum Dynamics on Modular Quantum Devices (CQD-MQD) under Award Number 2124511.
@@ V.1.1 Measurement of Kerr coefficient and frequency
The Kerr coefficient $K$, as used in the parametric oscillator Hamiltonian Eq. 2, is extracted by spectroscopy. A saturating probe drive $\omega_{pr}$ is applied to the SNAIL transmon operated with $\epsilon_{2}=0$. Varying the probe drive frequency and measuring the response via dispersive readout, results in the data shown in Fig. 7. The spectrum shows two clear dips, which correspond, from lower frequency to higher frequency, to the two-photon $gf/2$ transition and the $ge$ transition (the SNAIL levels are labeled following the atomic physics convention in increasing energy order $g,\;e,\;f,\;...$). Note that the measured $\omega_{ge}$ is a good approximation for $\omega_{a}$ used in the main text (Eq. 2). We then extract Kerr by using the relation $\omega_{ge}-\omega_{gf/2}=K$. Fitting two Gaussian peaks to the spectrum in Fig. 7, we find $K/2\pi=(528\pm 8)$ kHz. These measurements, when taken as a function of the biasing flux $\Phi$ of the superconducting loops, allow for a calibration of the SNAIL non-linear parameters $g_{3},g_{4}$. In Fig. 8 we show a flux scan of our sample with measured frequencies, Kerr nonlinearities, and the fit by the model presented in [17].
Figure 7: Continuous-wave spectroscopy measurement showing the readout response as a function of the probe tone frequency. From left to right, the pronounced dips in the signal show the $gf/2$ and $ge$ transitions of the SNAIL transmon. Those occur at $(\omega_{ge}-K)/2\pi$ and $\omega_{ge}/2\pi$ respectively. With $gf/2$, we refer to the two photon transition from ground state to second excited state. Fitting to the experimental data, we extract a Kerr value of $K/2\pi=(528\pm 8)$ kHz.
@@ V.1.2 Calibration of relative phase between squeezing drive and linear drive
Quantum coherent Rabi-like oscillations as a function of this phase $\phi$ are shown in Fig. 9 A [18]. For $\phi=0$, this drive is position-like and lifts the degeneracy between the two wells (see Fig. 2 C and Fig. 20 A). This energy difference induces the Rabi-like oscillation. For $\phi=90\degree$ (Fig. 9 A), the linear drive is momentum-like (the Hamiltonian term is $\propto\hat{p}$), thus not breaking the symmetry of the double-well (see Fig. 20 D).
Figure 8: Extraction of non-linearities from fit to frequency and Kerr parameter. A) Spectroscopic measurement of the oscillator’s frequency and B) its Kerr non-linearity as a function of flux as extracted from the $ge$ and $gf/2$ transitions (see Fig. 7). Solid lines are a simultaneous fit of frequency and non-linearity from the circuit model. We remark that several arrangements of parameters fit the data satisfactorily, producing a large correlation and errors (see [17] for the SNAIL model). In C) we show the third and fourth order non-linearities as extracted from the fit. The vertical dashed line represents the flux operating point for the experiments in the main text.
V.1.3 Calibration of squeezing (parametric) drive amplitude $\epsilon_{2}$ and linear drive amplitude $\epsilon_{1}$
In the experimental setup, the drive amplitudes are directly controlled by a digital to analog voltage converter. To calibrate the strength of the drive in MHz we measure time-resolved Rabi oscillations as a function of the digital control of the squeezing drive $\epsilon_{2}$. The experimental data is shown in Fig. 10 A. The oscillations of the observable $\hat{X}=(\ket{\alpha}\bra{\alpha}-\ket{-\alpha}\bra{-\alpha})/|\alpha|^{2}$ occur at a rate $\Omega_{\mathrm{cat}}(\epsilon_{2})\approx\Re(4\epsilon_{1}\alpha^{*})$ where $\alpha=\sqrt{\epsilon_{2}/K}$ and the approximation is valid for $|\alpha|>1$. Using, also, that for $\epsilon_{2}=0$ the oscillation has a frequency of $2\epsilon_{1}$ [18], just like for an ordinary transmon, we obtain $\epsilon_{1}$ in MHz completing the calibration of the drive amplitudes. To be clear, we can rewrite this relation as $\langle\hat{a}^{\dagger}\hat{a}\rangle=\epsilon_{2}/K=\Omega_{\mathrm{cat}}(\epsilon_{2})^{2}/16\epsilon_{1}^{2}$, where $\langle\hat{a}^{\dagger}\hat{a}\rangle$ is the average photon number of the coherent states. By extracting the Rabi rate $\Omega_{\mathrm{cat}}$ for each voltage of the digital control of $\epsilon_{2}$ and using the previously determined value of $\epsilon_{1}$, we find $\langle\hat{a}^{\dagger}\hat{a}\rangle$ as a function of the digital control of $\epsilon_{2}$. The data is shown in Fig. 10 B and shows a clear linear relationship between the applied voltage for the drive and the average photon number. The slope of the linear fit (together with the previously extracted value of Kerr) determines the proportionality constant between the digital to analog converter in volts and the drive amplitude $\epsilon_{2}$ in MHz as required by the Hamiltonian description.
@@
Figure 10: Calibration of drive strengths. A) Time-resolved quantum coherent Rabi-like oscillations as a function of squeezing amplitude. The squeezing amplitude is measured as the voltage of the digital controller. B) Photon number $\langle\hat{a}^{\dagger}\hat{a}\rangle$ as function of applied voltage for the digital control of squeezing drive $\epsilon_{2}$. The experimental data points are obtained from Fig. 10 A using $\langle\hat{a}^{\dagger}\hat{a}\rangle=\epsilon_{2}/K=\Omega_{\mathrm{cat}}^{2}/16\epsilon_{1}^{2}$. A linear fit allows us to convert the voltage set by the digital control of $\epsilon_{2}$ to the squeezing drive $\epsilon_{2}$ in MHz. In the main text, we only present data measured with up to $\epsilon_{2}/K=12$ (vertical dashed grey line).
@@ V.2.2 Steady state population
To investigate further the thermal properties of the system, we analysed the Lindbladian steady state $\rho_{\mathrm{ss}}$ in greater detail. To determine how well $\rho_{\mathrm{ss}}$ is described by a Boltzmann thermal state, we computed a collection of $\beta^{n,m}$ values using the detailed balance condition $p_{n}/p_{m}=\exp(-\beta^{n,m}(E_{n}-E_{m}))$, where $p_{n}$ are the eigenvalues of $\rho_{\mathrm{ss}}$ with eigenvectors closest to the eigenvectors $|E_{n}\rangle$ of $H_{\mathrm{eff}}$. We focused on pairs of neighbouring $p_{n}$ and $p_{m}$ values exceeding a small threshold ($10^{-6}$) and examined the distribution of $\beta^{n,m}$ values, as shown in Fig. 13 B and C. Additionally, we calculated the trace distance between $\rho_{\mathrm{ss}}$ and $\exp(-\beta_{\mathrm{avg}}H_{\mathrm{eff}})$, where $\beta_{\mathrm{avg}}$ is the average of all computed $\beta^{n,m}$ values. This is shown in Fig. 13 A. Overall, we found that the detailed balance condition is not well-satisfied for $\rho_{\mathrm{ss}}$, especially at the $1^{st}$ to $3^{rd}$ resonances (indicated by yellow lines), where the standard deviation of $\beta^{n,m}$ values is large, and their average appears to diverge. Notably, however, the condition holds moderately well between resonances and exceptionally well at the $4^{th}$ resonance, where the trace distance is approximately $10^{-3}$ and the standard deviation is around $10^{-2}$.
@@
Figure 13: Lindbladian simulation of detailed balance hypothesis at $\epsilon_{2}/K=7.7$, $\kappa/K=0.025$, $n_{\mathrm{th}}=0.05$ and $N=60$. A) Trace distance between Lindbladian steady state $\rho_{\mathrm{ss}}$ and Boltzmann distribution $\rho\sim\exp(-\beta_{\mathrm{avg}}H_{\mathrm{eff}})$ as a function of well’s asymmetry $\epsilon_{1}/K$. Note the near orthogonality at the 1st to 3rd resonance conditions (yellow lines) and the similarity at the 4th resonance condition (purple line). B) Average inverse temperature $\beta_{\mathrm{avg}}$ with error bars representing the standard deviation. C) Standard deviation of inverse temperature $\beta$. Note how the standard deviation becomes small at the 4th resonance condition (purple line).
@@ V.2.3 Control experiment for the relative phase dependence of the linear drive
The relative phase $\phi$ between the squeezing drive and linear drive is determined by searching for the phase of maximum Rabi rate, which occurs at $\phi=0$. The calibration measurement is shown in Fig. 9 A. This additional degree of freedom allows for several control experiments. To confirm that the resonant features we see are actually caused by the controlled symmetry breaking, and are not for example just power-dependent non-linear resonances [44, 45] or Stark shifting into resonance with spurious modes [46, 47], we repeat the experiment of Fig. 3 A for a symmetric double-well under an equally strong linear drive. To achieve this, we set the linear drive phase $\phi=90\degree$. The resulting measurement is shown in Fig. 9 B (see also Fig. 20 E). We note that the resonances are not present here, thereby proving that they are indeed a controlled effect from the symmetry breaking in the parametric oscilator.
@@
As discussed in the Sec. III, we have full control over the Rabi phase (see Fig. 9 A) which controls how the symmetric double-well is perturbed. A phase of $\phi=0\degree$ applies a drive that lifts the left well (see Fig. 20 A) and a phase of $\phi=90\degree$ does not break the $x$-symmetry (see Fig. 20 D). If a Rabi phase of $\phi=180\degree$ is applied, then the definitions of left and right well swap, meaning the right well becomes the shallower well and the left well becomes the deeper well. We can then run the same experiment again as in Fig. 3 A, meaning we initialize the system in the shallower (now right) well and then measure the population of the left well as a function of time and asymmetry. The resulting data (Fig. 9 C) shows the same increase in tunneling rate with larger asymmetry $\epsilon_{1}/K$ and resonances at specific values. The observation of the same physical phenomena confirms our understanding of the system and validates our explanations of the observed effects.
@@
Figure 15: Kerr correction. A) Reproduction of Fig. 3 A with a corrected Kerr to account for a $\approx 5\%$ uncertainty in the independent experimental value of $\epsilon_{1}/K$. That is, we use $(\epsilon_{1}/K)^{\prime}=1.05\times(\epsilon_{1}/K)$. There is only a minor improvement in the agreement between theory and experiment with respect to Fig. 3. B) Reproduction of Fig. 3 B.
@@
Figure 16: Kerr correction and additional visualization of resonance condition. A) Reproduction of Fig. 4 B with a $5\%$ Kerr adjustment like in Fig. 15 A. That is, we use $(\epsilon_{1}/K)^{\prime}=1.05\times(\epsilon_{1}/K)$ and $(\epsilon_{2}/K)^{\prime}=1.05\times(\epsilon_{2}/K)$. B) Lindbladian theory for comparison, now showing a slightly better agreement. C) Is another representation of the semiclassical approach to be compared with Fig. 4 A. In the latter we focus the attention on the phase quantization argument, while in this figure we focus the attention in the resonance condition between the levels in the asymmetric wells. Only the wells corresponding to the red circles are shown. The wells corresponding to the maroon circles are not shown.

### v2

@@ IV Relevance for chemistry
Figure 6: Lindbladian simulation of an ordinary double-well system. A) Dynamics simulated as a function of well-asymmetry $k_{1}/k_{4}$ and well-depth $k_{2}/k_{4}$. The color code marks the activation time (compare to Fig. 4 B). The green-black line shows the maximum activation time and the orange line marks the symmetric case ($k_{1}=0$). B) Comparison of the activation time for the ordinary symmetric ($k_{1}=0$) double-well and the (optimal) activation time along the green-black line in A as a function of well-depth $k_{2}/k_{4}$. C) Activation time as a function of asymmetry for $k_{2}/k_{4}=12.6$ (black dashed line in A), showing resonant tunneling (as in Fig. 3, see also [6]). Note that, as in the parametric oscillator Fig. 3, the linewidths alternate from broad to narrow.
The experimental observations presented here raise the question of whether they are present in other double-well systems, like the type of double-well involved in modeling chemical reactions. This may prove useful for investigations of problems relevant to chemistry (see, for example, [6] and references therein). To investigate this, we run a Lindbladian simulation for the chemical system with Hamiltonian $\hat{H}/\hbar=\hat{p}^{2}/2+k_{4}\hat{x}^{4}-k_{2}\hat{x}^{2}+k_{1}\hat{x}$, including single photon loss and gain with rate $\kappa/k_{4}=0.025$ and at temperature $n_{\mathrm{th}}=0.05$. These parameters lie in the ranges accessible by our system. Note that the dissipation model remains that of a high-quality factor linear oscillator, and we work under the hypothesis that this dissipation model captures the relevant phenomenology for applications in chemistry. With this, we will answer our question affirmatively: our experimental observations are predicted to be present in other double-well systems. On the one hand, the green-black line in Fig. 6 A depicts the maximum lifetime as a function of well-asymmetry and well-depth, showing that - just like in our experiment (Fig. 4 A, B) - a small asymmetry increases the lifetime significantly (see also Fig. 5 A, B and Fig. 6 B). On the other hand, Fig. 6 C depicts a horizontal linecut of Fig. 6 A and clearly reveals the same width alternation (broad-narrow-broad) as discovered and explained in our experiment (see Fig. 3). The finding of these two unexpected effects shows that our asymmetric Kerr parametric oscillator setup is already able to produce meaningful predictions for chemical quantum rate theory.
Double-well systems have been extensively studied in various contexts [33, 34], including in superconducting circuits [35, 36, 37, 38]. However, we speculate that the unique combination of precise real-time microwave control, complete tunability over Hamiltonian parameters, experimental stability, fast repetition rates, and high-fidelity readout in our setup explains why we observed effects not previously reported.
@@ V Conclusion
We also note the importance of this system for quantum computation since qubits can be encoded in the well state manifold [23, 17]. In this regard, two contributions of the present work deserve to be highlighted. The first one is the increase of the activation timescale $T$ by a fine control of the asymmetry. This leads to a reduction of bit-flip errors [43] with no extra hardware requirements. The second contribution is the demonstration of the operation of a highly asymmetric parametric oscillator in the quantum regime. We provide direct evidence that the static effective description is not compromised under strong linear drives, which are required for fast gates [23] and new implementations of hardware efficient readout schemes [44].
Our control over the well-asymmetry indicates that quantum parametric oscillators can implement analog quantum simulation of chemical reaction dynamics [39]. This allows for instance the analog simulation of proton tunneling, and e.g. the study of transfer reactions between the Guanine-Cytosine DNA base pairs, appears within reach of current Kerr parametric oscillator technology [39].
@@
ACCdA, RGC, and MS analyzed the data and modeled the experiment. ACCdA, RGC, MS, and MHD wrote the manuscript. PEV supported the data analysis. All authors contributed to the manuscript. RGC and MS conceptualized the experiment and collected data. NEF fabricated the device. BA, DGAC, PK, EG, and VSB validated the relevance and utility to chemistry and provided theory support. MHD and RGC supervised the research.
@@ Acknowledgements
We acknowledge useful discussions with Q. Su, A. Ding, B. Brock and M. Dykman. Remarks by V. Joshi and A. Koottandavida have improved the manuscript. This research was sponsored by the Army Research Office (ARO) under award number W911NF-23-1-0051. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office (ARO), or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. VSB, EG and MHD acknowledge partial support from the National Science Foundation Center for Quantum Dynamics on Modular Quantum Devices (CQD-MQD) under Award Number 2124511.
@@ V.1.1 Measurement of Kerr coefficient and frequency
The Kerr coefficient $K$, as used in the parametric oscillator Hamiltonian Eq. 2, is extracted by spectroscopy. A saturating probe drive $\omega_{pr}$ is applied to the SNAIL transmon operated with $\epsilon_{2}=0$. Varying the probe drive frequency and measuring the response via dispersive readout, results in the data shown in Fig. 7. The spectrum shows two clear dips, which correspond, from lower frequency to higher frequency, to the two-photon $gf/2$ transition and the $ge$ transition (the SNAIL levels are labeled following the atomic physics convention in increasing energy order $g,\;e,\;f$). Note that the measured $\omega_{ge}$ is a good approximation for $\omega_{a}$ used in the main text (Eq. 2). We then extract the Kerr frequency by using the relation $\omega_{ge}-\omega_{gf/2}=K$. Fitting two Gaussian peaks to the spectrum in Fig. 7, we find $K/2\pi=(528\pm 10)$ kHz. These measurements, when taken as a function of the biasing flux $\Phi$ of the superconducting loops, allow for a calibration of the SNAIL non-linear parameters $g_{3},g_{4}$. In Fig. 8 we show a flux scan of our sample with measured frequencies, Kerr nonlinearities, and the fit by the model presented in [19].
Figure 7: Continuous-wave spectroscopy measurement showing the readout response as a function of the probe tone frequency. From left to right, the pronounced dips in the signal show the $gf/2$ and $ge$ transitions of the SNAIL transmon. Those occur at $(\omega_{ge}-K)/2\pi$ and $\omega_{ge}/2\pi$ respectively. Fitting to the experimental data, we extract a Kerr value of $K/2\pi=(528\pm 10)$ kHz.
@@ V.1.2 Calibration of relative phase between squeezing drive and linear drive
Quantum coherent Rabi-like oscillations as a function of this phase $\phi$ are shown in Fig. 9 A [17]. For $\phi=0$, this drive is position-like and lifts the degeneracy between the two wells (see Fig. 2 C and Fig. 20 A). This energy difference induces the Rabi-like oscillation. For $\phi=90\degree$ (Fig. 9 A), the linear drive is momentum-like (the Hamiltonian term is $\propto\hat{p}$), thus not breaking the symmetry between the wells (see Fig. 20 D). The well-asymmetry is determined by $\epsilon_{1}\cos{\phi}$, and is therefore first-order insensitive to phase drifts of a few degrees around $\phi=0$, which is the precision of our calibration.
Figure 8: Extraction of non-linearities from fit to frequency and Kerr parameter. A) Spectroscopic measurement of the oscillator’s frequency and B) its Kerr non-linearity as a function of flux as extracted from the $ge$ and $gf/2$ transitions (see Fig. 7). Solid lines are a simultaneous fit of frequency and non-linearity from the SNAIL circuit model [19]. We remark that several arrangements of parameters fit the data satisfactorily, producing a large correlation and errors. In C) we show the third and fourth order non-linearities as extracted from the fit. The vertical dashed line represents the flux operating point for the experiments in the main text.
V.1.3 Calibration of the parametric squeezing drive amplitude $\epsilon_{2}$ and linear drive amplitude $\epsilon_{1}$
@@
[…]
Figure 10: Calibration of squeezing drive strengths. A) Time-resolved quantum coherent Rabi-like oscillations as a function of squeezing amplitude. The squeezing amplitude is controlled as the voltage of the digital controller. B) Photon number $\langle\hat{a}^{\dagger}\hat{a}\rangle$ as function of applied voltage for the digital control of squeezing drive $\epsilon_{2}$. The experimental data points are obtained from Fig. 10 A using $\langle\hat{a}^{\dagger}\hat{a}\rangle=\epsilon_{2}/K=\Omega_{\mathrm{Rabi}}^{2}/16\epsilon_{1}^{2}$, where $\epsilon_{1}$ is the asymmetry for $\phi=0$. A linear fit allows us to convert the voltage set by the digital control of $\epsilon_{2}$ to the squeezing drive $\epsilon_{2}$ in MHz. In the main text, we only present data measured with up to $\epsilon_{2}/K=12$ (vertical dashed grey line).
@@ V.2.2 Steady state population
To investigate the thermal properties of the system, we analysed the Lindbladian steady state $\hat{\rho}_{\mathrm{ss}}$. To determine how well $\hat{\rho}_{\mathrm{ss}}$ is described by a Boltzmann thermal state, we computed a collection of $\beta^{n,m}$ values using the detailed balance condition $p_{n}/p_{m}=\exp(-\beta^{n,m}(E_{n}-E_{m}))$, where $p_{n}$ are the eigenvalues of $\hat{\rho}_{\mathrm{ss}}$ with eigenvectors closest to the eigenvectors $|E_{n}\rangle$ of $\hat{H}_{\mathrm{eff}}$. We focused on pairs of neighbouring $p_{n}$ and $p_{m}$ values exceeding a small threshold ($10^{-6}$) and examined the distribution of $\beta^{n,m}$ values, as shown in Fig. 13 B and C. Additionally, we calculated the trace distance between $\hat{\rho}_{\mathrm{ss}}$ and $\exp(-\beta_{\mathrm{avg}}\hat{H}_{\mathrm{eff}})$, where $\beta_{\mathrm{avg}}$ is the average of all computed $\beta^{n,m}$ values. This is shown in Fig. 13 A. Overall, we found that the detailed balance condition is not well-satisfied for $\hat{\rho}_{\mathrm{ss}}$, especially at the $1^{\mathrm{st}}$ to $3^{\mathrm{rd}}$ resonances (indicated by yellow lines), where the standard deviation of $\beta^{n,m}$ values is large, and their average appears to diverge. Notably, however, the condition holds moderately well between resonances and exceptionally well at the $4^{\mathrm{th}}$ resonance, where the trace distance is approximately $10^{-3}$ and the standard deviation is around $10^{-2}$.
@@
Figure 13: Lindbladian simulation of detailed balance hypothesis at $\epsilon_{2}/K=7.7$, $\kappa/K=0.025$, $n_{\mathrm{th}}=0.05$ and $N=60$. A) Trace distance between Lindbladian steady state $\hat{\rho}_{\mathrm{ss}}$ and Boltzmann distribution $\hat{\rho}\sim\exp(-\beta_{\mathrm{avg}}\hat{H}_{\mathrm{eff}})$ as a function of well’s asymmetry $\epsilon_{1}/K$. Note the near orthogonality at the 1st to $3^{\mathrm{rd}}$ resonance conditions (yellow lines) and the similarity at the $4^{\mathrm{th}}$ resonance condition (purple line). B) Average inverse temperature $\beta_{\mathrm{avg}}$ with error bars representing the standard deviation. C) Standard deviation of inverse temperature $\beta$. Note how the standard deviation becomes small at the $4^{\mathrm{th}}$ resonance condition (purple line).
@@ V.2.3 Control experiment for the relative phase dependence of the linear drive
The relative phase $\phi$ between the squeezing drive and linear drive is determined by searching for the phase of maximum Rabi rate, which occurs at $\phi=0$. The calibration measurement is shown in Fig. 9 A. This additional degree of freedom allows for several control experiments. To confirm that the resonant features we measure are actually caused by the controlled symmetry breaking, and are not for example just power-dependent non-linear resonances [45, 46] or due to Stark shifting into resonance with spurious modes [47, 48], we repeat the experiment of Fig. 3 A for a symmetric double-well under an equally strong linear drive. To achieve this, we set the linear drive phase $\phi=90\degree$. The resulting measurement is shown in Fig. 9 B (see also Fig. 20 E). We note that the resonances are not present here, thereby proving that they are indeed a controlled effect from the symmetry breaking in the parametric oscillator.
@@
As discussed in the Sec. III, we have full control over the Rabi phase (see Fig. 9 A) which controls how the symmetric double-well is perturbed. A phase of $\phi=0\degree$ applies a drive that lifts the ”left” well (see Fig. 20 A) and a phase of $\phi=90\degree$ does not break the symmetry between the wells (see Fig. 20 D). If a Rabi phase of $\phi=180\degree$ is applied, then the definitions of left and right well swaps, meaning the right well becomes the shallower well and the left well becomes the deeper well. We can then run the same experiment again as in Fig. 3 A, meaning we initialize the system in the shallower (now right) well and then measure the population of the left well as a function of time and asymmetry. The resulting data (Fig. 9 C) shows the same increase in tunneling rate with larger asymmetry $\epsilon_{1}/K$ and resonances at specific values. The observation of the same physical phenomena confirms our understanding of the system and validates our explanations of the observed effects.
@@
Figure 15: Predictions based on fit procedure versus parameter-free theory prediction as in the main text. A) Reproduction of the data in Fig. 3 A with a corrected Kerr: we use $(\epsilon_{1}/K)^{\prime}=1.05\times(\epsilon_{1}/K)$. There is only a minor improvement in the agreement between theory and experiment with respect to Fig. 3. B) Reproduction of the data in Fig. 3 B.
@@
Figure 16: Kerr fit to data and alternative visualization of resonance condition. A) Reproduction of Fig. 4 B with a $5\%$ Kerr adjustment like in Fig. 15 A. That is, we use $(\epsilon_{1}/K)^{\prime}=1.05\times(\epsilon_{1}/K)$ and $(\epsilon_{2}/K)^{\prime}=1.05\times(\epsilon_{2}/K)$. B) Lindbladian theory for comparison, now showing a slightly better agreement. C) Is another representation of the semiclassical approach to be compared with Fig. 4 A. In the latter we focus the attention on the phase quantization argument, while in this figure we focus the attention on the resonance condition between the levels in the asymmetric wells. Only the wells corresponding to the red circles are shown. The wells corresponding to the maroon circles are not shown.